OpenStudy (anonymous):

Find the median. 12, 8, 5, 6, 1, 10, 13, 11

OpenStudy (anonymous):

DO I ADD OR MULTIPLY 10 AND 8 ?

Parth (parthkohli):

= (4th + 5th term) / 2 = (6 + 1)/2 = 3.5

Parth (parthkohli):

Note that two middle terms are done only when the number of data is even

OpenStudy (anonymous):

try this web site i really helped me http://www.mathsisfun.com/median.html